# A Comprehensive Study on Large Language Models for Mutation Testing

### v3

@@ 2.4. Evaluation Metrics
In this subsection, we present the evaluation metrics used in our study to assess the quality of the generated mutations comprehensively. We evaluate the mutations from multiple perspectives as before mentioned. Note that only relying on the mutation score is insufficient for a meaningful assessment. A higher mutation score does not necessarily indicate better quality mutations, as it only reflects the proportion of killed mutants rather than their effectiveness. However, we still report the mutation score of each mutation generation approach later in Section 3.
To better understand the metrics, we first illustrate the categories of mutations. Let all generated mutations be the set , where LLMs may generate syntactically incorrect mutations that can not pass compilation. Thus we denote the compilable mutations as the set . Note that the set contains two types of mutations that should be removed, the duplicate mutations (denoted as ) and equivalent mutations (denoted as ). Therefore, we have . The duplicate mutations refer to those identical to the original code or other mutations. In mutation testing, the ideal set of mutations to employ is from the set , excluding equivalent mutations that have no contributions to discovering the weakness of tests. However, determining the equivalent mutations set is undecidable, so we follow the common practice and use the set as a practical approximation to perform mutation testing (Diniz et al., 2021; Tian et al., 2022; Zhang et al., 2022; Wang et al., 2025).
@@ 2.4.1. RQ1: Effectiveness
To quantify this effectiveness, we employ multiple metrics that capture different aspects of behavioral similarity between mutations and real bugs. These metrics assess whether bug-triggering tests can detect LLM-generated mutations (i.e., RQ1.1), how closely mutations resemble real faults based on test case failures (i.e., RQ1.2), and the diversity of syntactic changes introduced by different mutation generation techniques (i.e., RQ1.3). The metrics of these sub-RQs are listed as bellow.
RQ1.1: Real Bug Detection Rate: Following the approach of Just et al. (Just et al., 2014b), we evaluate the effectiveness of mutation generation by measuring its ability to detect real bugs using developer-written tests. For a given real bug (which may span multiple lines), we apply a mutation generation approach to create a set of mutations on the fixed version of the program. If the test cases that kill a mutation also fail on the real bug, we consider that the mutation generation approach has successfully “detected” this real bug. To quantify this effectiveness, we compute the Real Bug Detection Rate, which represents the proportion of real bugs for which at least one generated mutation is detected by the same test cases that reveal the real bug. This metric provides insight into how well LLM-generated mutations align with actual software defects in a testing context.
@@
More concretely, let and be a mutation and its corresponding real bug, and let and be the set of their failing test cases respectively. The coefficient of is computed by the following formula:
@@
RQ1.2: Coupling Rate: Given a real bug, one mutant is considered coupling if the set of test cases that kill the mutant overlaps with the bug-triggering tests. Note that each bug we analyzed has at least one failing test to reveal the bug, called bug-triggering test. The Coupling Rate refers to the proportion of coupled mutants within the set of de-duplicated mutations (i.e., ). Note that Real Bug Detection Rate estimates the effectiveness via the proportion of detected bugs, which Coupling Rate estimates it via the proportion of mutations.
@@
RQ1.3: AST Node Diversity: We prefer the set of mutations with diverse syntactic forms, i.e., method invocations, field accesses, etc. Therefore, we parse the code before and after mutation, and check the diversity of newly introduced AST nodes. The approach with a greater variety of AST node types is considered superior, as it indicates a broader exploration of the syntactic space and enhances the likelihood of detecting weaknesses in the test suite.
RQ1.3: AST Edit Distance: The AST edit distance measures the structural difference between the original and mutated code by quantifying the number of tree transformation operations (i.e., insertions, deletions, and modifications) needed to convert one AST into another. This metric evaluates the extent of syntactic changes introduced by mutations. A higher edit distance suggests more significant syntactic changes.
@@ 2.4.2. RQ2: Validity
RQ2.1: Compilability Rate: the proportion of mutations that successfully compile, is computed as:
Because LLMs cannot guarantee the generation of entirely correct code and the compilation process can be time-consuming, we prefer the approach with a higher comparability rate.
@@
LLMs sometimes generate duplicate mutations, i.e., mutations that are exactly the same as the original code. While traditional approaches rarely generate duplicate mutations, LLMs can produce them due to their probabilistic text-generation mechanisms and learned patterns from large-scale training data.
@@
Equivalent mutations are semantically equivalent to the original program, which impacts the accuracy of the mutation score. We prefer the approach with a lower EMR.
@@ 2.4.3. RQ3: Efficiency
RQ3: Average Generation Time: measures the average time required to generate each mutation, providing a quantitative evaluation of the efficiency of each approach.
2.5. Mutation Generation
In this subsection, we discuss the setups for mutation generation approaches, especially the setups for LLM-based generation. We performed our experimental analysis on four rented cloud servers, each equipped with 2 NVIDIA GeForce RTX 3090 Ti, 168GB memory, and a 64 Core CPU Intel(R) Xeon(R) Platinum 8350C.
2.5.1. Mutation Generation via Traditional Approaches
For the well-established mutation generation approaches (i.e., PIT, Major, LEAM, and Bert), we activate all their mutation operators. Since we try to compare the similarity between the generated mutations and real bugs, we filter the mutations within the context of real-world bugs. Note that for all mutation generation approaches, we employ the same setting for the code context.
2.5.2. Mutation Generation via LLMs
Recently, LLMs have shown great potential in numerous code-related tasks, to better understand their ability in mutation testing, we also study LLMs for generating mutations.
LLMs operate based on prompts, which consist of natural language instructions, often accompanied by code snippets. LLMs are driven by prompts, which are natural language instructions maybe with code. To effectively leverage LLMs for mutation generation, we design a well-crafted prompt template, shown as Figure 1. We follow the best practices suggesting prompts should comprise four aspects, namely, Instruction, Context, Input Data, and Output Indicator (Guo et al., 2024; Ma et al., 2024). In the Instructions, we direct the LLMs to generate mutants for the target code element. In the Context, we clearly state that mutant is the concept of mutation testing, and additionally provide various sources of information to observe the impact on the performance of LLMs, including the whole Java method surrounding the target code element and few-shot examples sampled from real-world bugs from another benchmark. In the Input Data, we specify the code element to be mutated and the number of mutants to generate. In the Output Indicator, we specify the JSON file format for mutation outputs, facilitating further experiments. In addition, our prompts need few-shot examples, which should originate from real bugs enabling LLMs to learn how to mutate code from real bugs. This allows the model to understand the context and logic behind the changes, improving its ability to generate relevant and effective mutations.
To avoid the few-shot examples leaking information, beyond Defects4J and ConDefects, we employ another benchmark QuixBugs (Lin et al., 2017), which comprises 40 real-world Java bugs and is commonly used by the testing and debugging community (Zhu et al., 2023; Jiang et al., 2023; Xia and Zhang, 2022; Li et al., 2023b). Referring to the study by Deng et al. (Deng et al., 2024), LLMs perform best in generating valid code when using around six few-shot examples. Therefore, we sample six bugs from QuixBugs for few-shot learning in our prompt. To guarantee the diversity of the examples, we randomly select one bug from the dataset and assess whether its modification pattern is similar to the examples already collected. If it is different, we add it to our collection and continue, until we have collected all six examples, shown as Table 3.
| Correct Version | Buggy Version |
| n = (n & (n - 1)); | n = (n ^ (n - 1)); |
| while (!queue.isEmpty()) | while (true) |
| return depth==0; | return true; |
ArrayList r = new ArrayList();
r.add(first).addll(subset); to_add(r); |
to_add.addAll(subset); |
| c = bin_op.apply(b,a); | c = bin_op.apply(a,b); |
@@ 3. Evaluation Results
| Metric | PIT | Major | LEAM | BERT | GPT-3.5 | GPT-4o | GPT-4o-M | DS-236b | SC-16b | CL-13b |
| Mut. # | 340728 | 20831 | 36505 | 12946 | 31656 | 31945 | 34417 | 34741 | 28058 | 27614 |
| Mut. Score | 0.541 | 0.521 | 0.648 | 0.677 | 0.702 | 0.737 | 0.708 | 0.714 | 0.507 | 0.669 |
| Real Bug Detec. | 40.1% | 66.9% | 63.9% | 60.5% | 87.0% | 89.9% | 90.8% | 89.4% | 38.5% | 71.2% |
| Avg. Ochiai | 23.0% | 35.5% | 32.4% | 29.3% | 47.7% | 56.2% | 54.1% | 50.9% | 26.0% | 33.3% |
| Coupling Rate | 23.5% | 38.9% | 38.1% | 47.3% | 46.9% | 50.6% | 49.5% | 49.5% | 34.0% | 45.6% |
| Comp. Rate | 100% | 97.6% | 44.5% | 26.5% | 62.5% | 76.4% | 75.3% | 75.8% | 16.4% | 69.3% |
| Dup. Mut. Rate | 0.0% | 0.0% | 2.3% | 1.7% | 10.3% | 7.8% | 7.0% | 7.7% | 12.8% | 37.1% |
| A. G. Time (s) | 0.02 | 0.08 | 3.06 | 2.34 | 1.79 | 1.31 | 1.65 | 4.25 | 7.53 | 9.06 |
Note: The mutation generation approach achieving the best performance in terms of a metric is highlighted in grey background color. GPT-4o-M: GPT-4o-Mini. DS-236b: DeepSeek-Coder-V2-236b. SC-16b: StarChat--16b. CL-13b: CodeLlama-Instruct-13b.
@@ 3.1. RQ1: Performance on Effectiveness
Real Bug Real Bug Detection Rate: The row Real Bug Detec. of Table 4 presents the proportion of real bugs that can be detected across different mutation generation approaches. Among the traditional approaches, PIT achieves a detection rate of 40.1%, while Major, LEAM, and Bert perform significantly better with a detection rate of 66.9%, 63.9%, and 60.5%, respectively. For LLM-based approaches, the results are significantly more effective than those of traditional ones. GPT-4o-Mini achieves the highest detection rate at 90.8%, closely followed by GPT-4o and DeepSeek, with detection rates of 89.9% and 89.4%, respectively. Meanwhile, GPT-3.5 and CodeLlama achieve moderate detection rates of 87.0% and 71.2%, respectively.
| Project | Bug # | PIT | Major | LEAM | BERT | GPT-3.5 | GPT-4o | GPT-4o-M | DS-236b | SC-16b | CL-13b |
| Chart | 26 | 19 | 17 | 16 | 10 | 22 | 22 | 23 | 23 | 9 | 8 |
| Lang | 65 | 23 | 51 | 27 | 38 | 52 | 55 | 56 | 57 | 11 | 41 |
| Time | 27 | 12 | 23 | 18 | 15 | 25 | 24 | 26 | 22 | 10 | 17 |
| Math | 106 | 40 | 82 | 44 | 65 | 92 | 94 | 96 | 95 | 58 | 77 |
| Mockito | 38 | 14 | 28 | 27 | 5 | 30 | 29 | 36 | 34 | 16 | 21 |
| Closure | 133 | 38 | 92 | 122 | 103 | 130 | 131 | 131 | 132 | 79 | 114 |
| Cli | 39 | 6 | 11 | 27 | 29 | 37 | 39 | 39 | 39 | 2 | 34 |
| Codec | 18 | 6 | 11 | 12 | 11 | 14 | 16 | 17 | 17 | 1 | 11 |
| Csv | 16 | 7 | 13 | 13 | 7 | 15 | 16 | 16 | 16 | 2 | 13 |
| Gson | 18 | 6 | 9 | 11 | 12 | 13 | 16 | 16 | 15 | 1 | 10 |
| JacksonCore | 26 | 1 | 17 | 17 | 18 | 23 | 23 | 21 | 21 | 1 | 17 |
| Jsoup | 93 | 48 | 77 | 90 | 86 | 93 | 92 | 93 | 90 | 57 | 87 |
| ConDefects | 246 | 121 | 138 | 120 | 116 | 194 | 208 | 203 | 200 | 81 | 156 |
| ALL | 851 | 341 | 569 | 544 | 515 | 740 | 765 | 773 | 761 | 328 | 606 |
@@
To better understand the detection of the generation approaches, we analyze their project-wise detectability, shown as Table 5. For individual projects, LLM-based approaches consistently achieve top performance across multiple projects. Particularly, GPT-4o-Mini outperforms other approaches in several projects, including Chart, Time, Mockito, and Jsoup, where it detects the highest number of bugs. Similarly, DeepSeek and GPT-4o demonstrate strong performance, achieving the best results in projects such as Lang, Math, and Closure.
@@
Average Ochiai Coefficient: The row Avg. Ochiai of Table 4 presents the Ochiai Coefficient values across the approaches, which measures the semantic distance between mutations and real bugs. Among the traditional approaches, Major achieves an Ochiai coefficient of 35.5%, higher than the others. LEAM and BERT achieve coefficients of 32.4% and 29.3%, respectively, indicating that the best-performing rule-based approach, Major, outperforms both of them. The LLM-based approaches demonstrate significantly higher Ochiai coefficients, with GPT-4o achieving the highest value of 56.2%, followed closely by GPT-4o-Mini at 54.1%, and DeepSeek at 50.9%. Overall, GPT-4o and GPT-4o-Mini exhibit their superiority in generating mutations with smaller semantic distances with real bugs.
High Ochiai Bug Count: Following the study of Ojdanic et al. (Ojdanic et al., 2023), the mutation set with an Ochiai coefficient over 0.8 is considered to be of high semantic similarity with its corresponding real bug. For each project, we count the number of bugs that achieved a high Ochiai coefficient by each approach, shown as Table 6. The overall trend is similar with Real Bug Detectability, which indicates that the LLM-based appraoches, GPT-4o and GPT-4o-Mini, achieve the highest number of bugs across most projects. DeepSeek also exhibits strong performance, closely following GPT-4o in many cases. All traditional approaches exhibit significantly fewer high Ochiai bugs. Interestingly, among traditional approaches, Major outperforms LEAM and BERT across most projects.
| Project | Bug # | PIT | Major | LEAM | BERT | GPT-3.5 | GPT-4o | GPT-4o-M | DS-236b | SC-16b | CL-13b |
| Chart | 26 | 9 | 9 | 3 | 5 | 10 | 13 | 12 | 12 | 7 | 3 |
| Lang | 65 | 15 | 34 | 13 | 18 | 32 | 35 | 37 | 34 | 12 | 23 |
| Time | 27 | 8 | 14 | 8 | 4 | 8 | 12 | 12 | 11 | 4 | 7 |
| Math | 106 | 20 | 42 | 13 | 27 | 46 | 56 | 48 | 47 | 19 | 33 |
| Mockito | 38 | 1 | 12 | 5 | 3 | 14 | 13 | 17 | 14 | 3 | 3 |
| Closure | 133 | 12 | 25 | 40 | 14 | 38 | 69 | 56 | 45 | 21 | 26 |
| Cli | 39 | 3 | 3 | 13 | 8 | 18 | 22 | 23 | 20 | 4 | 10 |
| Codec | 18 | 3 | 1 | 3 | 3 | 2 | 8 | 6 | 7 | 3 | 2 |
| Csv | 16 | 4 | 2 | 1 | 2 | 4 | 10 | 10 | 8 | 3 | 1 |
| Gson | 18 | 2 | 1 | 4 | 3 | 2 | 8 | 7 | 6 | 3 | 2 |
| JacksonCore | 26 | 2 | 3 | 9 | 8 | 11 | 14 | 12 | 11 | 5 | 6 |
| Jsoup | 93 | 16 | 6 | 32 | 20 | 46 | 39 | 39 | 41 | 25 | 19 |
| ConDefects | 246 | 101 | 150 | 132 | 134 | 175 | 179 | 181 | 177 | 112 | 148 |
| ALL | 851 | 196 | 302 | 276 | 249 | 406 | 478 | 460 | 433 | 221 | 283 |

### v4

@@ 2.4. Evaluation Metrics
In this subsection, we present the evaluation metrics used in our study to assess the quality of the generated mutations comprehensively. We evaluate the mutations from multiple perspectives, as mentioned before. Note that only relying on the mutation score is insufficient for a meaningful assessment. A higher mutation score does not necessarily indicate better quality mutations, as it only reflects the proportion of killed mutants rather than their effectiveness. However, we still report the mutation score of each mutation generation approach later in Section 3.
To better understand the metrics, we first illustrate the categories of mutations. Let all generated mutations be the set , where LLMs may generate syntactically incorrect mutations that can not pass compilation. Thus, we denote the compilable mutations as the set . Note that the set contains two types of mutations that should be removed, the duplicate mutations (denoted as ) and equivalent mutations (denoted as ). Therefore, we have . The duplicate mutations refer to those identical to the original code or other mutations. In mutation testing, the ideal set of mutations to employ is from the set , excluding equivalent mutations that have no contribution to discovering the weakness of tests. However, determining the equivalent mutations set is undecidable, so we follow the common practice and use the set as a practical approximation to perform mutation testing (Diniz et al., 2021; Tian et al., 2022; Zhang et al., 2022; Wang et al., 2025). Therefore, let denote the set of killed mutants among the compilable and non-duplicate ones (i.e., ). The mutation score is then computed as follows:
Note that traditional mutation generation approaches, such as Major and PIT, produce very few duplicate mutations (i.e., is 0). Hence, early studies (Kintis et al., 2017) defined equivalent mutations as those semantically equivalent to the original code, which may implicitly include , which is an empty set. However, LLM-based approaches tend to generate a large number of non-mutated duplicates. Moreover, while duplicate mutations can be easily filtered, identifying equivalent mutations remains challenging (Tian et al., 2024). Therefore, in this paper, we refine the definition of equivalent mutations to be those that are syntactically different but semantically equivalent to the original code.
@@ 2.4.1. RQ1: Effectiveness
To quantify this effectiveness, we employ multiple metrics that capture different aspects of behavioral similarity between mutations and real bugs. These metrics assess whether bug-triggering tests can detect LLM-generated mutations (i.e., RQ1.1), how closely mutations resemble real faults based on test case failures (i.e., RQ1.2), and the diversity of syntactic changes introduced by different mutation generation techniques (i.e., RQ1.3). The metrics of these sub-RQs are listed below.
RQ1.1: Real Bug Detection Rate: Following the approach of Just et al. (Just et al., 2014c), we evaluate the effectiveness of mutation generation by measuring its ability to detect real bugs using developer-written tests. For a given real bug (which may span multiple lines), we apply a mutation generation approach to create a set of mutations on the fixed version of the program. If the test cases that kill a mutation also fail on the real bug, we consider that the mutation generation approach has successfully “detected” this real bug. To quantify this effectiveness, we compute the Real Bug Detection Rate, which represents the proportion of real bugs for which at least one generated mutation is detected by the same test cases that reveal the real bug. This metric provides insight into how well LLM-generated mutations align with actual software defects in a testing context.
@@
More concretely, let and be a mutation and its corresponding real bug, and let and be the set of their failing test cases, respectively. The coefficient of is computed by the following formula:
@@
RQ1.2: Coupling Rate: Given a real bug, one mutant is considered coupling if the set of test cases that kill the mutant overlaps with the bug-triggering tests. Note that each bug we analyzed has at least one failing test to reveal the bug, called bug-triggering test. The Coupling Rate refers to the proportion of coupled mutants within the set of de-duplicated mutations (i.e., ). Note that Real Bug Detection Rate and Coupling Rate measure effectiveness from different perspectives. Real Bug Detection Rate evaluates effectiveness at the bug level, i.e., the proportion of bugs that can be revealed by at least one mutant. In contrast, Coupling Rate evaluates effectiveness at the mutation level, i.e., the proportion of mutants whose failing tests overlap with bug-triggering tests. For instance, a mutation approach may generate a large number of high-quality mutants for only a few specific bugs, but few mutants for others. This can result in a high Coupling Rate (since the generated mutants are strongly aligned with the bug-triggering tests of those few bugs) but a low Real Bug Detection Rate (since many other bugs remain undetected). Conversely, another approach may generate mutants that cover a wide range of bugs, thereby achieving a high Real Bug Detection Rate. However, if the mutants for each bug exhibit only weak coupling with the bug-triggering tests, the overall Coupling Rate will remain low. Therefore, these two metrics are complementary, with each reflecting only one aspect of mutation effectiveness.
@@
RQ1.3: AST Node Diversity: We prefer the set of mutations with diverse syntactic forms, i.e., method invocations, field accesses, etc. Therefore, we parse the code before and after mutation, and check the diversity of newly introduced AST nodes. In other words, we identify the unique AST node types introduced by each approach’s generated mutants, relative to the set of node types present in the original program. The approach with a greater variety of AST node types is considered superior, as it indicates a broader exploration of the syntactic space and enhances the likelihood of detecting weaknesses in the test suite.
RQ1.3: AST Edit Distance: The AST edit distance measures the structural difference between the original and mutated code by quantifying the number of tree transformation operations (i.e., insertions, deletions, and modifications) needed to convert one AST into another. This metric evaluates the extent of syntactic changes introduced by mutations. A higher edit distance suggests more significant syntactic changes. The AST Edit Distance metric is well-suited for structured code, as it captures syntactic differences at the tree level rather than merely comparing token sequences, as in BLEU and similar metrics. However, it still has inherent limitations, such as its inability to fully reflect the impact of mutations on program behavior. We further discuss these limitations later in Section 5.
@@ 2.4.2. RQ2: Validity
RQ2.1: Compilability Rate: the proportion of mutations that successfully compile, which is computed as:
Because LLMs cannot guarantee the generation of entirely correct code and the compilation process can be time-consuming, we prefer the approach with a higher compilability rate.
@@
LLMs sometimes generate duplicate mutations, i.e., mutations that are either identical to the original code or identical to each other. While traditional approaches rarely generate duplicate mutations, LLMs can produce them due to their probabilistic text-generation mechanisms and learned patterns from large-scale training data. After eliminating duplicate mutations, all remaining mutations are guaranteed to be syntactically distinct both from the original code and from each other.
@@
Equivalent mutations are semantically equivalent to the original program, which impacts the accuracy of the mutation score. We prefer the approach with a lower EMR. Note that while equivalent mutations are also considered semantically duplicate, they are not easily identifiable. Therefore, we distinguish them from other duplicate mutations, which can be filtered syntactically.
@@ 2.4.3. RQ3: Efficiency
RQ3.1: Average Generation Time: measures the average time required to generate each mutation, providing a quantitative evaluation of the efficiency of each approach.
RQ3.2: Average Token: measures the average number of consumed tokens per mutation, reflecting both the financial cost (due to token-based pricing) and the computational time cost (since longer token sequences typically require more processing time). Since input and output token costs differ across models, we calculate them separately.
2.5. Mutation Generation Setups
In this subsection, we discuss the setups for mutation generation approaches, especially the setups for LLM-based generation.
For the well-established mutation generation approaches (i.e., PIT, Major, LEAM, and BERT), we activate all their mutation operators. Since we try to compare the similarity between the generated mutations and real bugs, we filter the mutations within the context of real-world bugs. Note that we supply the same code snippets as context across all mutation generation approaches to ensure a fair comparison.
For all LLM-based approaches, we use the default settings of each model (e.g., token limits and temperature). For models accessible via Web API (e.g., GPT series and DeepSeek models), we rely on their official APIs. For the remaining open-source LLMs, we perform inference on four rented cloud servers, each equipped with dual NVIDIA GeForce RTX 3090 Ti GPUs, 168 GB of memory, and a 64-core Intel(R) Xeon(R) Platinum 8350C processor.
@@ 3. Evaluation Results
| Name | #Mut. | Mut. S. | R. B. Detec. | Avg. Ochiai | Coup. Rate | Comp. Rate | Dup. Rate | Time |
| PIT | 340728 | 0.541 | 40.1% | 23.0% | 23.5% | 100% | 0.0% | 0.02 |
| Major | 20831 | 0.521 | 66.9% | 35.5% | 38.9% | 97% | 0.0% | 0.08 |
| LEAM | 36505 | 0.648 | 63.9% | 32.4% | 38.1% | 44.5% | 2.3% | 3.06 |
| BERT | 12946 | 0.677 | 60.5% | 29.3% | 47.3% | 26.5% | 1.7% | 2.34 |
| BF (GPT-4o) | 13068 | 0.762 | 58.6% | 17.7% | 47.9% | 74.1% | 7.5% | 5.35 |
| LLMph (GPT-4o) | 20043 | 0.754 | 68.5% | 24.1% | 51.7% | 67.8% | 7.1% | 3.41 |
| LLMut (GPT-3.5) | 31656 | 0.702 | 87.0% | 47.7% | 46.9% | 62.5% | 10.3% | 1.79 |
| LLMut (GPT-4o) | 31945 | 0.737 | 89.9% | 56.2% | 50.6% | 76.4% | 7.8% | 1.31 |
| LLMut (GPT-4o-M) | 34417 | 0.708 | 90.8% | 54.1% | 49.5% | 75.3% | 7.0% | 1.65 |
| LLMut (DS-761b) | 33596 | 0.749 | 91.1% | 59.8% | 50.8% | 77.6% | 7.5% | 2.57 |
| LLMut (DS-236b) | 34741 | 0.714 | 89.4% | 50.9% | 49.5% | 75.8% | 7.7% | 4.25 |
| LLMut (SC-16b) | 28058 | 0.507 | 38.5% | 21.7% | 34.0% | 16.4% | 12.8% | 7.53 |
| LLMut (CL-13b) | 27614 | 0.669 | 71.2% | 33.3% | 45.6% | 69.3% | 37.1% | 9.06 |
Note: The mutation generation approach achieving the best performance in terms of a metric is highlighted in a grey background color.
BF: BugFarm. LLMph: LLMorpheus. LLMut: Our new prompt.
GPT-4o-M: GPT-4o-Mini. DS-761b: DeepSeek-V3-761b. DS-236b: DeepSeek-Coder-V2-236b. SC-16b: StarChat--16b. CL-13b: CodeLlama-Instruct-13b.
@@ 3.1. RQ1: Performance on Effectiveness
Real Bug Detection Rate: The column Real Bug Detec. of Table 4 presents the proportion of real bugs that can be detected across different mutation generation approaches.
Among the traditional approaches, PIT achieves a detection rate of 40.1%, while Major, LEAM, and BERT perform significantly better with a detection rate of 66.9%, 63.9%, and 60.5%, respectively.
For LLM-based approaches, the results are significantly more effective than those of traditional ones. The existing LLM-based approaches, BugFarm and LLMorpheus, achieve a rate of 58.6% and 68.5%, respectively. For the LLMs based on our prompting, DeepSeek-V3-761b achieves the highest detection rate at 91.1%, closely followed by GPT-4o-Mini, GPT-4o and DeepSeek-Coder-V2-236b, with detection rates of 90.8%, 89.9%, and 89.4%, respectively. Meanwhile, GPT-3.5 and CodeLlama achieve moderate detection rates of 87.0% and 71.2%, respectively.
| Proj. | Chart | Lang | Time | Math | Moc. | Clo. | Cli | Co. | Gs. | Csv | Ja. | Js. | ConD. | All |
| #Bug | 26 | 65 | 27 | 106 | 38 | 133 | 39 | 18 | 16 | 18 | 26 | 93 | 246 | 851 |
| PIT | 19 | 23 | 12 | 40 | 14 | 38 | 6 | 6 | 7 | 6 | 1 | 48 | 121 | 341 |
| Major | 17 | 51 | 23 | 82 | 28 | 92 | 11 | 11 | 13 | 9 | 17 | 77 | 138 | 569 |
| LEAM | 16 | 27 | 18 | 44 | 27 | 122 | 27 | 12 | 13 | 11 | 17 | 90 | 120 | 544 |
| BERT | 10 | 38 | 15 | 65 | 5 | 103 | 29 | 11 | 7 | 12 | 18 | 86 | 116 | 515 |
| BF (GPT-4o) | 14 | 37 | 15 | 57 | 28 | 93 | 23 | 6 | 8 | 11 | 7 | 83 | 117 | 499 |
| LLMph (GPT-4o) | 17 | 51 | 20 | 61 | 31 | 121 | 30 | 12 | 8 | 12 | 9 | 87 | 124 | 583 |
| LLMut (GPT-3.5) | 22 | 52 | 25 | 92 | 30 | 130 | 37 | 14 | 15 | 13 | 23 | 93 | 194 | 740 |
| LLMut (GPT-4o) | 22 | 55 | 24 | 94 | 29 | 131 | 39 | 16 | 16 | 16 | 23 | 92 | 208 | 765 |
| LLMut (GPT-4o-M) | 23 | 56 | 26 | 96 | 36 | 131 | 39 | 17 | 16 | 16 | 21 | 93 | 203 | 773 |
| LLMut (DS-761b) | 24 | 61 | 27 | 101 | 35 | 131 | 39 | 13 | 16 | 16 | 25 | 92 | 195 | 775 |
| LLMut (DS-236b) | 23 | 57 | 22 | 95 | 34 | 132 | 39 | 17 | 16 | 15 | 21 | 90 | 200 | 761 |
| LLMut (SC-16b) | 9 | 11 | 10 | 58 | 16 | 79 | 2 | 1 | 2 | 1 | 1 | 57 | 81 | 328 |
| LLMut (CL-13b) | 8 | 41 | 17 | 77 | 21 | 114 | 34 | 11 | 13 | 10 | 17 | 87 | 156 | 606 |
@@
BF: BugFarm. LLMph: LLMorpheus. LLMut: our prompt. Moc.: Mockito. Clo.: Closure. Co.: Codec. Gs.: Gson. Ja.: JacksonCore. Js.: Jsoup. ConD: ConDefects.
To better understand the detection of the generation approaches, we analyze their project-wise detectability, shown as Table 5. For individual projects, LLM-based approaches consistently achieve top performance across multiple projects. Notably, DeepSeek-V3-761b detects the largest number of bugs in most Defects4J projects and achieves the highest overall bug detection count. Similarly, DeepSeek-Coder-V2-236b, GPT-4o, and GPT-4o-Mini demonstrate strong performance, achieving the best results in projects such as Mockio, Csv, and ConDefects.
@@
Average Ochiai Coefficient: The column Avg. Ochiai of Table 4 presents the Ochiai Coefficient values across the approaches, which measures the semantic distance between mutations and real bugs. Among the traditional approaches, Major achieves an Ochiai coefficient of 35.5%, higher than the others. LEAM and BERT achieve coefficients of 32.4% and 29.3%, respectively, indicating that the best-performing rule-based approach, Major, outperforms both of them. Existing other LLM-based approaches achieve an overall lower Ochiai coefficient of 17.7% and 24.1%, respectively. The LLM-based approaches using our prompt achieve notably higher Ochiai coefficients. DeepSeek-V3-761b attains the highest value at 59.8%, followed by GPT-4o at 56.2% and GPT-4o-Mini at 54.1%. The LLMs based on our prompt outperform other existing methods, not only traditional approaches but also existing LLM-based ones, showing their superiority in generating mutations that are semantically closer to real bugs.
High Ochiai Bug Count: Following the study of Ojdanic et al. (Ojdanic et al., 2023), the mutation set with an Ochiai coefficient over 0.8 is considered to be of high semantic similarity with its corresponding real bug. For each project, we count the number of bugs that achieved a high Ochiai coefficient by each approach, shown as Table 6. The overall trend is similar to Real Bug Detectability, which indicates that the LLM DeepSeek-V3-761b, based on our prompt, achieves the highest number of bugs across most projects. The closed-source LLMs GPT-4o and GPT-4o-Mini also exhibit strong performance, closely following DeepSeek-V3-761b in many cases. Compared with our approach, all traditional approaches exhibit significantly fewer high Ochiai bugs. Interestingly, among traditional approaches, Major outperforms LEAM and BERT across most projects.
| Proj. | Chart | Lang | Time | Math | Moc. | Clo. | Cli | Co. | Gs. | Csv | Ja. | Js. | ConD. | All |
| #Bug | 26 | 65 | 27 | 106 | 38 | 133 | 39 | 18 | 16 | 18 | 26 | 93 | 246 | 851 |
| PIT | 9 | 15 | 8 | 20 | 1 | 12 | 3 | 3 | 4 | 2 | 2 | 16 | 101 | 196 |
| Major | 9 | 34 | 14 | 42 | 12 | 25 | 3 | 1 | 2 | 1 | 3 | 6 | 150 | 302 |
| LEAM | 3 | 13 | 8 | 13 | 5 | 40 | 13 | 3 | 1 | 4 | 9 | 32 | 132 | 276 |
| BERT | 5 | 18 | 4 | 27 | 3 | 14 | 8 | 3 | 2 | 3 | 8 | 20 | 134 | 249 |
| BF (GPT-4o) | 9 | 15 | 3 | 13 | 1 | 13 | 3 | 2 | 1 | 4 | 2 | 14 | 71 | 151 |
| LLMph (GPT-4o) | 6 | 25 | 7 | 24 | 6 | 22 | 9 | 2 | 3 | 3 | 5 | 12 | 81 | 205 |
| LLMut (GPT-3.5) | 10 | 32 | 8 | 46 | 14 | 38 | 18 | 2 | 4 | 2 | 11 | 46 | 175 | 406 |
| LLMut (GPT-4o) | 13 | 35 | 12 | 56 | 13 | 69 | 22 | 8 | 10 | 8 | 14 | 39 | 179 | 478 |
| LLMut (GPT-4o-M) | 12 | 37 | 12 | 48 | 17 | 56 | 23 | 6 | 10 | 7 | 12 | 39 | 181 | 460 |
| LLMut (DS-761b) | 13 | 40 | 15 | 62 | 21 | 79 | 27 | 8 | 10 | 12 | 17 | 59 | 146 | 509 |
| LLMut (DS-236b) | 12 | 34 | 11 | 47 | 14 | 45 | 20 | 7 | 8 | 6 | 11 | 41 | 177 | 433 |
| LLMut (SC-16b) | 7 | 12 | 4 | 19 | 3 | 21 | 4 | 3 | 3 | 3 | 5 | 25 | 112 | 221 |
| LLMut (CL-13b) | 3 | 23 | 7 | 33 | 3 | 26 | 10 | 2 | 1 | 2 | 6 | 19 | 148 | 283 |